Two rockets of rest length L approach each other with equal speeds of 0.4c. How
long would one rocket appear to the other?
1
Expert's answer
2020-04-13T10:23:44-0400

Since both rockets move at the speed comparable to the speed of light, one rocket will notice that the length of the other will decrease following Lorentz transformation:


"L_c=L\sqrt{1-v^2/c^2}=\\=L\sqrt{1-(0.4c)^2/c^2}=0.917L."